7. Graph the equation in Desmos or Geogebra. Then answer the following questions. A person throws a ball straight up in the air. The height of a ball, h, in meters, can be modelled by h=-4.9t2+10.78t+1.071, where t is the time in seconds since the ball was thrown.
a)What is the maximum height the ball can reach?

b)How tall is the person that throws the ball up in the air?

c)If the thrower also wants to catch the ball, what time do you think that will occur at?

Answer by htmentor(1343)

h(t) = -4.9t2+10.78t+1.071
The time is takes to reach maximum height is given by t = -b/2a = -10.78/(-2*4.9) = 1.1 s
a) h(1.1) = 7 m
b) The height of the person is the height at t = 0, h(0) = 1.071
c) The roundtrip time to return to the original height is 2*1.1 = 2.2 s